OpenStudy (anonymous):

Kristina reaches into a bag of candy that has 12 diet candies and 6 candies with sugar. She randomly pulls out one, puts it back, and picks out another. What is the probability that both candies she pulled out have sugar?

OpenStudy (anonymous):

Probability is the same for each candy. It isn't affected by how many candies you pick before or after that candy, as long as you put the candy back each time. So the probability for the first candy = probability for the second candy = number of sugar candies/total number of candies

OpenStudy (anonymous):

*100%, I think, if you want to express it as a percentage

OpenStudy (anonymous):

Thank ya! :)

OpenStudy (anonymous):

:D

OpenStudy (accessdenied):

The probability of each event is the same (they don't affect each other, they're independent), but we're looking for the probability of "A and B", which is the product of the probabilities 6/16 = 3/8 3/8 * 3/8 = 9/64

OpenStudy (anonymous):

^ Good job AD, I didn't pick up on that.